OpenStudy (anonymous):

I will fan and medal! Miguel and Maria are washing the windows in their home. Together, they can complete the task in 1.5 hours. If Maria can wash windows twice as fast as Miguel, how many minutes would it take her to wash them by herself? In your answer, include the equation you used to solve the problem.

OpenStudy (anonymous):

i might be able to help

OpenStudy (anonymous):

so x = 1.5

OpenStudy (anonymous):

maria = 1/2x (twice as fast)

OpenStudy (anonymous):

your solving for maria but...

OpenStudy (anonymous):

i might be wrong but the equation is 1/2x * x = 1.5 and then you solve for x

OpenStudy (anonymous):

if m = rate of Miguel, then 2m is the rate of maria. Given they work together and fish in 1.5 hrs, we have 1.5 (m + 2m) = 1 m = 2/9. Therefor, the rate of Maria is 2(2/9) = 4/9, thus it will take her 1/ (4/9) = 9/4 hrs to complete the task and 9/4 hours = 2 hrs 15 min